### v1

@@ Abstract
Artificial engineering often relies on scarce materials, possessing ecological, social, and political risks. In contrast, nature achieves advanced functionalities using only abundant elements and water. Despite this simple insight, incorporating nature-inspired devices made from readily available resources into industrial applications remains a formidable challenge. Here, we present a sustainable ’blue battery’—a water-based energy storage device that mimics natural processes and reduces dependence on scarce materials. This battery harnesses the distinctive electrical properties of water confined within nanometer-scale pores, enabling construction from abundant, non-toxic materials using scalable nanofabrication techniques. It operates without unwanted side reactions, maintains nearly 100% efficiency after 60,000 charge-discharge cycles, and supports a voltage window of up to 1.65 V while achieving competitive power and energy density. The core component, a van der Waals clay nanostructure, uses water as the sole electrolyte with enhanced activity due to interactions with its 2D surface. This water-driven technology offers a sustainable platform for energy systems, with potential applications from small-scale electronics to global energy infrastructure—even in extreme environments like Mars. Our work advances nature-inspired energy solutions, opening new possibilities for the sustainable engineering of blue devices.
1 Main
In contrast to man-made engineering approaches, nature builds its functionalities using abundant ’primitive’ chemical elements and water, prioritizing nanoscale structuring over the optimization of chemical composition [1, 2, 3]. This strategy, which ensures scalability and growth critical for survival [4], presents a blueprint for engineering innovations aimed at sustainability. By mimicking nature, we can reduce dependence on scarce materials and develop water-based systems—hereafter referred to as ’blue devices’—which hold the potential to meet rising demands in areas such as sustainable batteries [5], neurointerfaces [6], and neuromorphic computing [7]. However, a thorough understanding of water’s atomic and molecular behavior at interfaces, a topic of growing interest in recent years [8, 9, 10, 11, 12, 13], remains essential for advancing these technologies.
Recent studies have revealed unusual properties of water confined in pores ranging from a few angstroms to several nanometers, or ultraconfinement [14, 15, 16]. Under these extreme conditions, water exhibits behavior distinct from bulk water, challenging existing theoretical models [17]. Experiments have revealed unique effects such as dielectric polarization anisotropy and protonic ’superconductivity,’ both linked to altered proton dynamics in ultraconfined water [18, 19, 20, 21]. These discoveries open exciting possibilities for blue devices, particularly in improving aqueous batteries and supercapacitors. These devices offer safer, more cost-effective alternatives to traditional technologies that rely on metal oxides and concentrated electrolytes [22, 23], and are crucial for stabilizing renewable energy networks powered by solar and wind. While previous efforts in nanofluidic systems have demonstrated impressive performance in energy applications [24, 25], scaling them up for practical use remains a significant challenge [26, 27]. Utilizing ultraconfined water in nature-inspired designs with abundant materials could unlock new scaling-up possibilities while addressing environmental and sustainability concerns.
In this work, inspired by nature, we introduce a sustainable ’blue battery.’ This water-based device utilizes pure water confined within extremely narrow 1-nm-high slit pores as an electrolyte, where proton activity is significantly enhanced compared to bulk water. We demonstrate that few-molecule-high water channels facilitate the separation and accumulation of protons and hydroxides within the electrodes’ electrical double layers (EDLs). To ensure practical applicability, the device is constructed exclusively from abundant materials—clay, graphene, and water—reducing reliance on scarce resources in energy storage. The device achieves up to 100% coulombic efficiency, with a projected equivalent lifespan of 150 years without significant degradation. It operates within a voltage window of 1.65 V, offering competitive power and energy densities. By focusing on precise structural design rather than rare chemical elements, our ’blue battery’ provides a scalable and sustainable platform for various future technologies. These include safe energy storage systems, electrolyzers, bio-computer interfaces, and neuromorphic devices, where ultraconfined water within solids serves as both an electrolyte and a separator.
1.1 Topologically connected channels as blue device concept
Unlike previous aqueous electrochemical systems with geometrically disjointed electrodes (Fig. 1, a and b), we introduce a topologically connected membrane-electrode unit (Fig. 1c), which mimics natural systems such as the interstitial fluid in biological organisms (Fig. 1d), providing structural continuity within tissue and organ boundaries [28]. Consider the simplest aqueous capacitor, the ’Leyden jar’ [29], which consists of two metal electrodes in contact with water (Fig. 1a). In this device, electric energy is stored in a thin, nanometer-scale interface (blue glow) with a low energy density due to a negligible surface-to-volume ratio. To improve this, the classic battery or supercapacitor design (Fig. 1b) increases the surface area of the electrodes to enhance energy density but also increases an electrolyte concentration to secure high current density, thereby promoting chemical complexity. In contrast, we propose transforming the disjointed electrodes into a unified, topologically connected network of electrical double layers (EDLs) without involving bulk water (Fig. 1c). This design leverages the unique properties of ultraconfined water without increasing chemical complexity, maintaining sustainability without compromising performance.
To design our blue device, we created an electrode-membrane-electrode unit out of abundant 2D materials loaded with ultraconfined water within 1-nm-high slit channels (Fig. 1e). In the channels the walls are so close that the double layers of water from opposite sides overlap, eliminating bulk water and leaving only interfacial water with affected self-dissociation-recombination balance [17]. Such water exhibits modified electrodynamic properties and an increased activity of H+ and OH- ions. Additionally, the EDLs of the electrodes connect topologically through the EDLs of the separator, creating a continuous network across the device parts. This design mimics nature by focusing on space structuring rather than chemical diversity. As demonstrated below, the device achieves exceptional cycle life and competitive performance. Water access into the aligned slit pores (Fig. 1f) is ensured by the use of hydrophilic clay, whose properties require a separate analysis at the microscopic level.
1.2 Where water meets Earth: clays as natural 2D materials
Clays occur on all Earth’s continents (Fig. 2a) and also on Mars [30]. On the microscopic level, they have fine crystalline structure (Fig. 2b) [31]. The most common types of clay—kaolinite, illite, and smectite (Fig. 2c)—consist of 1-nm-thick crystals made of silica and alumina planes, interchanging differently depending on the clay type to form flat flakes. These flakes are held together by vdW interactions, forming periodic heterostructures up to micrometers in size (SI Figs. S1-S4). Natural clays are chemically simple and composed of silicon, oxygen, and aluminum atoms but may also include water-soluble potassium, sodium, or calcium, which can be reversibly substituted by, e.g. hydrogen [31, 32].
Unlike synthetic vdW heterostructures [34, 35], clays offer several advantages: low price, scalability, widespread availability, natural affinity to water, safety, and biocompatibility. Their organized slit channels (Fig. 2b) facilitate water and proton transport [32], allowing for the movement and relative separation of positive and negative ions under natural conditions or an external electric field. Besides intercellular spaces (Fig. 1d), the clay interlayer pores are similar to small wood pores or nerve cell structures [36, 37], making them suitable for nature-mimicking engineering.
Our X-ray diffraction (XRD) analysis showed that exposing dried clays to saturated water vapor or liquid water results in interlayer expansion or swelling, evidenced by diffraction peak shifts (Fig. 2e and SI Figs. S5 and S6). This process and the natural clay structure allow the creation of periodic slit channels filled with water (Fig. 2d). Notably, this procedure is much less expensive and time-consuming than the artificial vdW channels design [38] and has almost infinite materials sourcing capabilities. We found that smectite, in particular, has an equilibrium pore height of 1 nm, consistent with previous reports [39, 40]. Coincidentally, these dimensions correspond to the peak maximum of proton conductivity observed in water confined in a nanoporous medium and within artificial slit pores [18, 20, 21].
To test the electrical properties of clays, we first carefully cleaned the samples and then hydrated them in saturated water vapor (see Methods for sample preparation). This process ensured the absence of contamination and foreign ionic impurities. Dry clays acted as good insulators, while wet clays conducted ions with the assistance of water. We observed significant ionic conductivity in all wet clays, in contrast to both bulk water and dry clays, which exhibited high resistivity (1 Mcm, Fig. 2f). Smectite, in particular, demonstrated the highest direct-current (DC) conductivity of 0.6 S/m, comparable to that found in synthetic slits and ceramics [21, 20], and equivalent to the conductivity of a 0.1-M NaCl solution in bulk water or ionic liquids [41]. Notably, this conductivity is six orders of magnitude greater than that of bulk water, which is 0.610-7 S/m [17] (see blue dashed line).
@@ 1.2 Where water meets Earth: clays as natural 2D materials
The enhanced ionic transport in hydrated clay can be understood via dynamic interactions between the water and the clay under conditions of ultraconfinement. The natural substitution of Si4+ and Al3+ atoms in the clay’s crystal structure by Al3+ and Mg2+, respectively, creates an excess negative surface charge [31]. In nature, this excess surface charge is balanced by adsorbed alkaline cations to maintain electro-neutrality [31]. In cleaned clays, this role is played by H+ (excess protons) of water, whose adsorption and desorption increase the conductivity of confined water by disrupting its self-dissociation equilibrium [42] and raising the concentration and mobility of H3O+ and OH- (see inset in Fig. 2f). This effect is most pronounced in pores around 1.5 nm (approximately 2 times the Bjerrum length, nm), aligning with experimental data [20, 21] and predictions from the ionic model of water [17]. Notably, excess protons move via a relay-race proton-hopping mechanism, with water molecules acting as mediators, resulting in significantly higher mobility compared to other ions [17].
This scenario is supported by the negative surface charge we detected on colloidal nanoparticles of clays in distilled water (SI Table S1). The interaction between clay and water enables ultrahigh ion conduction within the 1-nm slit channels, transforming confined water into a partially ionic liquid with H+ and OH- ions as charge carriers. This effect, observed in other nanoconfinement systems [18, 20, 21], provides opportunities for nature-inspired engineering of blue devices. The use of clay, which processing does not require any extra chemicals unlike other 2D materials, ensures low cost and natural hydration of the system.
1.3 Energy storage in confined water: blue battery
To utilize the unique properties of ultraconfined water, including its inherent safety, giant in-plane polarizability [21], and enhanced conductance from surface-water interactions [20], we designed an electricity storage unit (blue battery). Using the nanopore vacuum filtration technique (Fig. 3a; see Methods), we sequentially and controllably self-assembled layer by layer the mixed graphene-clay vdW heterostructures to form two composite electrodes divided by a pure clay separator. The result was a scalable, free-standing membrane-electrode unit (MEU) in the form of nanocomposite film approximately 200 m thick (Fig. 3b, c), featuring a three-layer structure (Fig. 3d), and having high stability and flexibility, typical for vdW heterostructures [38].
SEM EDX mapping (Fig. 3e, f) confirmed the presence of only clay (O, Al, Si) and graphene (C) elements (see SI Figs. S9, S10 for additional data). Electron microscopy showed that within the layers (Fig. 3g, h), material flakes are aligned to form parallel channels, which intersect at the edges of the domains (Fig. 3i). Moreover, no voids are detected at the interface between the domains of the separator and the electrodes. The channels were filled with water via capillary condensation from saturated water vapor, with mass gain used to control the process. This design promotes ultraconfined water intercalation into the electrodes and the separator while eliminating bulk-water voids. The structure aligns with the topological design shown in Fig. 1c. Notably, this scalable system requires no additional chemicals and is highly reproducible using standard laboratory procedures.
@@ 1.3 Energy storage in confined water: blue battery
To test the device, two graphite current collectors were attached to both sides of the MEU (Fig. 4a). The cell underwent standard electrochemical tests within a 2.1 V voltage window (see Methods) to evaluate charge-discharge cycles (Fig. 4b), current-voltage characteristics (Fig. 4c), and estimate capacitance, coulombic efficiency, and energy efficiency. Despite its simple design (Fig. 4d) and use of pure water without additives, the cell showed impressive parameters and durability with no visible degradation after 60,000 cycles (Fig. 4e). This cycle life translates to over 150 years of operation counting one charge-discharge cycle per day, e.g., in the case of use with solar panels. It retains and releases energy with up to 100% coulombic and 60% energy efficiency, due to the absence of side reactions. The initial ”annealing” period was assigned to the electric-field-stimulated wetting of nanopores.
@@
Moreover, we found that the device has an electrolysis threshold of 1.65 V (Fig. 4f, and Fig. S11), higher than the standard 1.23-V limit for bulk water at neutral conditions, which yields higher capacitance, and energy density than that in bulk water systems. This is likely due to the combined effects of high water purity and the ultraconfinement of water, which alters its electrodynamics, particularly conductivity (Fig. 2f) and polarizability [18, 19, 21]. Our system achieves a capacitance of around 40 F/g and an energy density of around 10 Wh/kg of electrode material, comparable to that of commercial supercapacitors [43]. Both parameters have the potential for further improvement through optimization.
Note that the device’s efficiency and energy density significantly exceed those of biological systems such as photosynthesis, which has an efficiency below 1% [44]. Thus, our device mimics, but also surpasses nature. It traps and transports protons and hydroxides (proton holes) of the nanometer-thick water-solid interface within the cell over macroscopic distances in an external electric field and stores them in the same-size nanopores in the electrodes (see Fig. 1e). This effect is universal and can be used in other types of electrochemical systems, such as electrolyzers, batteries, and fuel cells, as well as nature-signaling mimicking systems and neuromorphic devices.
@@ 2 Discussion
We have demonstrated a scalable platform for energy applications that leverages the anomalous properties of ultraconfined water: an enhanced proton conductivity [20] and heterogeneous polarizability [21]. The system is topologically non-equivalent to the standard electrochemical systems design (Fig. 1, a to c) and mimics nature’s use of ultraconfinement and space (Fig. 1d). Our device, made from only abundant materials: graphene, clay, and water confined within 1-nm slit pores, serves as a rechargeable electricity retention unit with a projected lifespan of over 150 years (Fig. 4e). This makes it ideal for supplementing solar or wind generators to avoid grid instability, which is an urgent problem of a fast-growing renewable energy network. The exceptional durability and infinite materials sourcing suggest it could be a reliable solution for balancing electricity generation-consumption time lags or providing fast-frequency control in the grid [45]. A performance comparison with current batteries and supercapacitors indicates that our system could offer superior long-term stability that suggests specific niche applications (SI Table S3).
Note that while using rare elements in energy devices can enhance reaction speed, energy density, efficiency, and operational ranges (voltage or temperature), our nature-mimicking approach suggests advantages in terms of cost, environmental safety, geopolitical stability, biocompatibility, scalability, structural stability, and durability. The non-toxic nature of clay [46] makes our platform also suitable for a range of bio-applications, including battery pills [47], bio-interfaces [6], neuromorphic computing [7], and even for use in Mars colonization, where clays and water are available on-site [48], whereas ultraconfined water does not freeze down to -70 ∘C [49].
In conclusion, our blue-battery platform represents a significant advance in applying nature-inspired principles to sustainable nanoengineering of energy systems. By combining 2D materials science with abundant materials and simple fabrication, we achieve high performance, safety, material independence, and sustainability. Our technology aims to drive further research and seek sustainable industrial applications, promoting a more efficient, environmentally friendly, and safer energy future.

### v3

@@ Abstract
Nanometer-scale solid-state confinement has been shown to change water’s structure and dynamics, offering new horizons in energy storage. However, most current materials operate at the micrometer scale, missing the interfacial effects that occur at three orders of magnitude smaller dimensions. Here, we report a scalable energy storage device that uses ultraconfined water as its sole electrolyte, unlocking the advantages of nanoscale confinement. We use the polarizability and proton ’superconductivity’ of water confined in few-molecular-diameters clay channels to build an all-water supercapacitor. The device fabricated from reconstructed clay, graphene, and water by a sustainable self-assembly process, operates at voltages up to 1.65 V, has competitive power and energy density, and maintains near 100% Coulombic efficiency over 60,000 charge-discharge cycles. These results demonstrate the application of unique properties of ultraconfined water for sustainable energy storage and provide a benchmark for a class of novel ultraconfined water-based energy systems, or ’blue devices’.
1 Introduction
The ultraconfinement of liquids within solid-state nanopores has gained exceptional attention over the past decade due to its critical role in emerging technologies, including osmotic power generation, ultrafiltration, and molecular sensing [1, 2, 3, 4, 5]. A liquid becomes ultraconfined when the electrical double layers (EDLs) from opposing surfaces overlap. In these conditions, the liquid exhibits drastically different—and sometimes counterintuitive—electrodynamic properties compared to bulk liquid [6]. The deviations in the rheological, mechanical, and electrodynamic properties of ultraconfined liquids in the pores of a few nanometers are also evident from the myriads of natural interfacial phenomena [7, 8]. Despite the extensive fundamental studies [9, 10, 11], the electrodynamic properties of ultraconfined liquids remain essentially unexploited because the nanochannels fabrication and measurements at the nanoscale are challenging.
Recent studies have shown dielectric anomalies in ultraconfined pure water, revealing structural heterogeneity and unusual molecular dynamics. Specifically, the dielectric constant anisotropy and the effect of proton ’superconductivity’ of ultraconfined water have been experimentally observed [12, 13, 14, 15] and supported by simulations [16]. These effects challenge our understanding of water’s bulk properties and suggest intriguing applications. Notably, they open up possibilities for sustainable and safer electrochemical energy storage systems, such as supercapacitors, by utilizing the topology and strong confinement instead of increasing chemical complexity. This approach mirrors nature’s strategy of emphasizing shape and structure over increasing chemical species to develop functionality [17], and yet to be explored.
Unlike phenomena in traditional aqueous supercapacitors, which utilize water as a solvent for concentrated electrolytes, the ultraconfinement affects the intrinsic electrodynamic properties of water without changing its atomic composition. This phenomenon has been understood via the surface-water interaction, which affects its self-ionization followed by charge separation [18]. The altered electrodynamics of ultraconfined water has been observed in various nanoporous systems, including that of hydrophilic and hydrophobic properties [12, 14, 19, 15]. This means that the phenomenon has a universal nature on a broader range of functional materials beyond the polymers traditionally used to enhance proton transport [20]. Although an attempt was made to create an all-water electrochemical supercapacitor using nanodiamond ceramics [19], the limited scalability of this and other nanoporous platforms has restricted the application of ultraconfined water in large-scale electricity storage.
Here, we introduce a scalable energy storage platform, called the ’blue device’, which utilizes electrical double-layer capacitance of ultraconfined pure water. By harnessing the enhanced proton conductivity of up to 0.6 S/m within 1-nm water channels in nanoengineered naturally abundant clays, we achieve energy and power densities competitive with conventional supercapacitors, hybrid capacitors, and some aqueous batteries. Operating at 1.65 V with nearly 100% Coulombic efficiency and a projected lifespan of 150 years, the device offers a highly sustainable, safe, and eco-friendly solution for bulk electricity storage.
2 Results
2.1 Topologically connected channels and blue device concept
The simplest aqueous capacitor is the Leyden jar [21] (Fig. 1a). It consists of two non-porous electrodes in contact with water. The electric energy is stored in a thin nm-high interface (EDL). Though sustainable, the device has negligible for practical applications energy density due to a small surface-to-volume ratio and low concentration of charge carriers able to travel between the electrodes. In contrast, the modern aqueous supercapacitor has large-surface-area electrodes immersed in a concentrated electrolyte or an ionic liquid (Fig. 1b). It suggests a relatively high power and energy density yet has chemical complexity. The latter increases the price and raises sustainability issues.
To address these problems, we propose transforming the disjointed electrodes into a unified, topologically connected network of nanopores lasting continuously between and within the electrodes through the nanoporous separator without a break (Fig. 1c). This design leverages the unique properties of water ultraconfined in nanochannels. Particularly, benefiting from the proton ’superconductivity’ and anomalous polarizability without increasing chemical complexity. The design is topologically nonequivalent to the previous two and maintains sustainability like the Leyden jar, but without compromising performance compared to standard supercapacitors.
A corresponding to the sketch in Fig. 1c device membrane-electrode unit (MEU) (Fig. 2a) can be assembled by using abundant clay and graphene nanosheets. This design suggests a continuum of water-filled nanochannels (Fig. 2b) between periodic solid-state sheets.
Unlike synthetic vdW heterostructures [22, 23, 24], natural clay nanosheets offer several advantages. They are low-cost, scalable, widely available, water-affine, safe, and biocompatible. Slit-like channels between the clay sheets (Fig. 2b) enable efficient proton transport [25]. In some ways, this design resembles the small pores in wood or the interstitial spaces between cells [26, 27], making clay sheets ideal for nature-inspired nanoengineering.
2.2 Clays as natural 2D materials
To identify the best clay materials for the blue device, we tested the electrical and hydrophilic properties of several abundant clays, including kaolinite, illite, and smectite (see SI Fig. S1). These clays naturally form small crystals (Fig. 3a) that can grow to micrometer sizes (SI Figs. S3-S6). Found across all continents (see SI Fig. S2) and even on Mars [28], clays consist of 1-nm-thick crystal sheets made of silica and alumina planes (Fig. 3b) [29]. The sheets are weakly bound, allowing interlayer water penetration and swelling, and can be reconstructed to form large-scale heterostructures.
To ensure the material’s purity and to exclude the dissolution of foreign species into the water channels, all the raw samples were carefully cleaned before use by multiple centrifugation and dialysis, followed by energy dispersive X-ray elemental analysis (see Methods). The water pH and conductivity were measured before and after the measurements. Finally, water was introduced to the channels by adsorption from saturated water vapor excluding foreign species presence.
The hydrophilic properties of purified clays were tested using X-ray diffraction (XRD) by assessing interlayer spacing between periodic crystal sheets. Among the studied clays, smectite exhibited the highest swelling and was selected for further experiments. The wet smectite has an interlayer distance expansion of about 1 nm compared to dry one (Fig. 3c and SI Figs. S5, S6), consistent with previous studies [30, 31]. Thus, nearly 50% of the wet clay volume consists of water-filled pores, with the remaining portion being the solid matrix.
Dielectric properties of clays were compared for dry and wet conditions by electrical conductivity spectrum measurements. Dry clays were found to be good insulators (Fig. 3d), while cleaned and hydrated samples exhibited ionic conductivity far exceeding that of bulk water (see blue dashed line for reference). Smectite, in particular, showed the highest direct-current conductivity at 0.6 S/m—comparable to polymer electrolyte membranes like Nafion [32] and bulk ionic liquids [33]. Thus, it was assigned as the best for the design of the blue device.
@@ 2.2 Clays as natural 2D materials
Enhanced proton transport in hydrated clay can be understood by considering dynamic interactions between water and the clay-crystal walls under ultraconfinement. In the crystal structure, natural substitution of Si4+ and Al3+ with Al3+ and Mg2+ [29] creates an excess negative surface charge (SI Table S1). In nature, this charge is balanced by adsorbed alkaline cations [29], but in cleaned clays, hydrogen ions (H+) from water take over this role. This increases conductivity by disrupting water’s self-dissociation equilibrium [34], raising the concentration and mobility of H3O+ and OH-. Thus, the ultraconfinement of water in clays transforms the former into a partially ionic liquid with hydronium and hydroxyl ions as charge carriers without increasing chemical complexity.
The maximum surface conductivity of ultraconfined water is expected in pores around twice the Bjerrum length nm—which marks the balance between electrostatic interactions and thermal energy (kBT). Indeed, experimentally, peak proton conductivity has been reported in water confined within nanopores of around 2 nm in various materials [12, 14, 15].
2.3 Energy storage in 1-nm-high water channels
To fabricate the blue device, we used a nanopore filtration technique (Fig. 4a). We assembled layer-by-layer graphene-clay composite electrodes separated by a pure clay membrane (Fig. 4b) from corresponding colloidal suspensions (see Methods). We produced dozens of membrane-electrode units (MEUs) with varying layer thicknesses and graphene-clay compositions (see SI Table S2). All MEUs (Fig. 4c) demonstrated excellent reproducibility, mechanical stability, and flexibility.
Fig. 4d shows a cross-sectional image of a MEU and Figs. 4e, f demonstrate corresponding EDX mapping. This analysis confirms the presence of only native clay elements (O, Al, Si) and graphene (C), as prescribed following the preliminary cleaning procedure (see SI Figs. S9, S10 for further data). Electron microscopy (SEM and STEM) shows (Fig. 4, g to i,) that reconstructed clay sheets align to form parallel channels, intersecting at domain edges and providing a pathway for molecular and ionic transport.
To test the device, the nanopores were loaded with water by exposure of MEU to the saturated water vapor. Two graphite current collectors were applied to both sides of the MEU (Fig. 5a) to avoid direct contact of metal connectors with water. The cell underwent standard electrochemical tests within a voltage window up to 2.1 V (see Methods). The charge-discharge cycles (Fig. 5b) and current-voltage characteristics (Fig. 5c) were measured, and the capacitance, coulombic efficiency, and energy efficiency were estimated. Fig. 5d shows the MEU charge-discharge mechanism by separating water’s hydronium and hydroxyl ions.
@@ 2.3 Energy storage in 1-nm-high water channels
The cell showed impressive parameters and durability with no visible degradation after 60,000 cycles (Fig. 5e). This cycle life translates to over 150 years of operation counting one charge-discharge cycle per day, e.g., in the case of use with solar panels. It retains and releases energy with up to 100% Coulombic and 60% energy efficiency, due to the absence of side reactions. The initial ”annealing” period was assigned to the electric-field-stimulated wetting of nanopores.
@@
We found that the device has an electrolysis threshold of 1.65 V (Fig. 5f, and Fig. S11), higher than the standard 1.23-V limit for bulk water at neutral conditions, which yields higher capacitance, and energy density than that in bulk water systems. This is likely due to the combined effects of high water purity and the ultraconfinement of water, which alters its electrodynamics, particularly conductivity (Fig. 3f) and polarizability [12, 13, 15]. Our system achieves a capacitance of around 40 F/g and an energy density of around 10 Wh/kg of electrode material, competitive with that of commercial supercapacitors, hybrid capacitors, and some batteries [35]. Both parameters have the potential for further optimization.
@@ 3 Discussion
We have introduced a scalable and safe platform—the ”blue device”—for electric energy storage in ultraconfined water, leveraging its unique electrodynamic properties. Unlike conventional aqueous supercapacitors, our system takes a fundamentally different approach by enabling topological crosslinking of interfacial water layers across electrodes and separators. This design, achieved using abundant materials and simple fabrication methods, offers high performance, safety, and long-term sustainability without relying on scarce resources. It exemplifies nature-inspired principles in nanoengineering.